Write the point slope form of a line that passes through (6,1) and is perpendicular to a line with a slope of -3
Answer by Alan3354(69443)

Write the point slope form of a line that passes through (6,1) and is perpendicular to a line with a slope of -3
---------------
Perpendicular --> the slope, m, is the negative inverse, = 1/3
---------
Use y-y1 = m*(x-x1) where (x1,y1) is the point.
y-1 = (1/3)*(x-6)
---
To get point-slope form, solve for y.